OpenStudy (anonymous):

how can I solve these my self???? 4d+2=6?

OpenStudy (anonymous):

transpose 2 to right hand side 4d = 6-2 4d = 4 divide both sides by 4 4d/4 = 4/4 d= 1

OpenStudy (anonymous):

got it or not roseanne i hope u must be satisfied with my solution

OpenStudy (anonymous):

Hey, Roseanne, how r u? Weren't u posting much more difficult questions than this a while back (my apologies if I mixed you up with someone else)?

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

it was my pleasure roseanne